Calculate maximum no of radiations emitted of infrared region when an electron falls from 5th excited state to ground state 

Grade:10

1 Answers

Shruti
5 years ago
Since 5th excited state means n=6.                              Therefore, number of radiations are given by the relation = n(n-1)÷2.                                                                       =(6×5)÷2=15 is the correct answer.